OpenStudy (anonymous):

Kayla bought 3 notebooks and 3 pens for $6.00. Camilo bought 2 notebooks and 4 pens for $5.00. How much did 1 notebook cost?

OpenStudy (anonymous):

[3N +3P=6] 4 ----> 12N +12P =24 -[2N +4P=5] 3----->-(6N +12P = 15) ------------------ 6N = 9 N=9/6 = $1.50

OpenStudy (anonymous):

thanks booo (: